# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v3

@@ 3.3. Galois Closed QC Codes and Their Duals
Let
$R_{q^{2},m}^{\ell}\cong\bigoplus_{j=1}^{t}\left(\dfrac{\mathbb{F}_{q^{2}}[x]}{\langle g_{j}\rangle}\oplus\dfrac{\mathbb{F}_{q^{2}}[x]}{\langle g^{*}_{j}\rangle}\right)^{\ell}\oplus\bigg(\bigoplus_{i=1}^{s}\dfrac{\mathbb{F}_{q^{2}}[x]}{\langle f_{i}\rangle}\bigg)^{\ell},$
where $(g_{j},g_{j}^{*})$ are pairs of reciprocal irreducible factors, for $1\leq j\leq t$, and $f_{i}$ are self-reciprocal irreducible factors, for $1\leq i\leq s$, with $f_{s}=x-1$. Let $\mathcal{C}_{u}$ be the $m^{u}\ell$-length QC code whose constituent codes (40) and (41) are all Galois closed. Additionally, assume that the constituent codes (41) are HSO. Then, we have an infinite family $\{\mathcal{C}_{u}\}_{u\geq 2}$ of HSO $q^{2}$-ary QC codes satisfying the square-root-like lower bound for their minimum distances.
4. Constructions of QECCs from QC Codes
Given the construction of $\ell$-QC codes and results seen in [cem, lingsole], along with the collection of quantum constructions in Subsection 2.4, we present in this section a construction of quantum codes with good parameters.
@@ 4. Constructions of QECCs from QC Codes
$\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{7}-1\rangle}\right)^{3}$ $\cong$ $\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{3}+x+1\rangle}\right)^{3}\oplus\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{3}+x^{2}+1\rangle}\right)^{3}\oplus\left(\frac{\mathbb{F}_{4}[x]}{\langle x-1\rangle}\right)^{3}$
@@
For any element $\gamma\in\mathbb{F}_{4}(\alpha)\setminus\mathbb{F}_{4}$, let $C_{11}^{\prime}=\left\langle(\gamma,\gamma,\gamma)\right\rangle\subset\mathbb{F}_{4}^{3}(\alpha)$ , $C_{11}^{\prime\prime}=(C_{11}^{\prime})^{\perp}\subset\mathbb{F}_{4}^{3}(\alpha^{3})$, and $C_{11}=\mathbb{F}_{4}^{3}$ be the constituent codes, whose parameters are $3,1,3]_{64}$, $3,2,2]_{64}$, and $3,3,1]_{4}$, respectively. Notice that $C_{11}^{\prime},C_{11}^{\prime\prime}$, and $C_{11}$ are Galois closed by Proposition 3.12, and $C_{11}$ is trivially HDC. Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC-code which by Theorem 3.13 is also Galois-closed. The minimum distance of the cyclic code associated to $\mathcal{C}$ and all its associated cyclic subcodes are described below
@@
By Theorem 2.9 (See Remark 2), the minimum distance of the $\ell$-QC code $\mathcal{C}$ is lower-bounded by
(60) $d(\mathcal{C})\geq\min\{7,7,8\}=7,$
and $\mathcal{C}$ is a $21,12,\geq 7]_{4}-$EDC $\ell$-QC code by [lingsole, Theorem 4.2]. As seen in Corollary 2.11, from $\mathcal{C}$ it is possible to obtain a $21,3,d^{\perp}\geq 7]]_{4}-$quantum code $Q$. According to [table], the best known quaternary QECC for $n=21$ and $k=3$ has minimum distance $7.$
By the Lemma 2.12, it is still possible to construct more good QECCs from the explicit code given above which are listed below:
@@
Ex. 4.1 and Lemma 2.12 - item (2) $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$ $\geq 6$
Code Table [table] 4-5 5-6 5-7 6-8 7-8
@@
Ex. 4.1 and Lemma 2.12 - item (1) $\geq 7$ $\geq 7$ $\geq 7$ $\geq 7$ $\geq 7$
Code Table [table] 7-10 8-11 8-11 9-11 9-12
Although there is no significant improvement in the minimum distances for any of the above parameters, most of our codes achieve values that are within one or two of the best-known lower bounds. It is also worth noting that our example exhibits lower bounds of the minimum distance, so it may be possible to obtain codes with larger minimum distances.
@@
Still working with the decomposition of $R_{4,7}^{3}$ provided in Example 4.1, assume the same constituent codes $C_{11}^{\prime}$ and $C_{11}^{{}^{\prime\prime}}$, and let $C_{11}=\langle(0,1,\beta),(1,0,\beta^{2})\rangle$ be the last constituent code, where $\beta\in\mathbb{F}_{4}$ so that $\beta^{2}=\beta+1$, whose parameters are $3,2,2]_{4}$, respectively. Observe that $C_{11}$ is EDC.
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC-code. By Theorem 3.7, we have an infinity family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of EDC QC codes whose their parameters are
(61) $\left[3\cdot 7^{u},3\left(\frac{7^{u}+1}{2}\right),\geq 2\cdot 7^{u-1}\right]_{4}.$
From Corollary 2.11, it is possible to get the infinity family of quantum codes $\{Q_{u}\}_{u\geq 1}$ with parameters $\left[\left[3\cdot 7^{u},3,\geq 2\cdot 7^{u-1}\right]\right]_{4}$.
Example 4.3.
Let $q=2$, $m=7$, $\ell=8$, and $\alpha$ be a primitive $7$-th root of unity. Assume the following decomposition of $R_{2,7}^{8}=\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$. Thus
@@
Let $C_{11}^{\prime}$ be a $8,3,6]_{8}$ MDS code (its construction can be checked in [table]), $C_{11}^{{}^{\prime\prime}}=C_{11}^{{}^{\prime}\perp}$ a $8,5,4]_{8}$ (also MDS) code, and $C_{11}$ be a EDC $8,6,2]_{2}-$code with basis
(62) $B$ $=$ $\left\{1,0,0,1,0,0,0,0),(0,1,0,1,0,0,0,0),(0,0,1,1,0,0,0,0),(0,0,0,0,1,0,0,1),(0,0,0,0,0,1,0,1),\right$.
$\left.(0,0,0,0,0,0,1,1)\right\}$
From the minimum distances of the associated cyclic subcodes provided in the Example 4.1, together with the minimum distances of the constituent codes described above, the minimum distance of $\mathcal{C}$ in $\mathbb{F}_{2}^{56}$ is $d\geq d_{GO}=14.$ Moreover, by Lemma 2.1, the dimension of $\mathcal{C}$ is 30, namely, $\mathcal{C}$ is a $8$-QC EDC $56,30,\geq 14]_{2}-$code and, based on Corollary 2.11, it is possible to construct a QECC with parameters $56,4,\geq 14]]_{2}$. Notice that in [table], the best $56,4]]_{2}-$QECC has minimum distance 13.
By the Lemma 2.12, we list more good quantum codes from this code. We have an improvement on some lower bounds from this explicit example: the best minimum distance for $55,5]]_{2}$ can be improved to 13–18 and $57,4]]_{2}$ can be improved to 14–19.
$51,9]]_{2} [[$$52,8]]_{2} [[$$53,7]]_{2} [[$$4,6]]_{2} [[55$$5,5]]_{2}$
Ex. 4.3 and Lemma 2.12 - item (2) $\geq 9$ $\geq 10$ $\geq 11$ $\geq 12$ $\geq 13$
Code Table [table] 11–15 11–16 11–17 12–17 12–18
$57,4]]_{2} [[$$58,4]]_{2} [[$$59,4]]_{2} [[$$60,4]]_{2} [[$$61,4]]_{2}$
Ex. 4.3 and Lemma 2.12 - item (1) $\geq 14$ $\geq 14$ $\geq 14$ $\geq 14$ $\geq 14$
Code Table [table] 13–19 14–19 14–19 14–20 14–20
5. Conclusion
@@
In general, describing the parameters of infinity families of codes is not an easy task. In this paper, we have presented families of QC codes over $\mathbb{F}_{q}$, making explicit their dimension and a lower bound to their minimum distances. In addition, based on their constituent code decomposition, we also explored their dualities according to Euclidean and Hermitian inner products. We have also addressed when this lower bound to minimum distance is greater than the square-root-like bound, which there are only a few constructions in the literature achieving this goal. Finally, some quantum codes have been reported, which are comparable with those ones depicted on the Code Table [table].
@@
For future work, we aim to gain a deeper understanding of how using MDS codes as constituent codes affects the minimum distance of the respective QC code. It is also important to explore whether the GO-bound can be improved, since doing so would lead to better lower-bound estimates for the minimum distances of both our QC codes and their corresponding quantum codes.
6. acknowledgements
@@
The authors would like to thank the Institute for Computational and Experimental Research in Mathematics (ICERM) for providing funding and space for us to work in a mathematically rich environment during our week-long visit in Summer 2025. The first two authors would also like to thank Simons Laufer Mathematical Sciences Institute (SLMath) for providing funding for post-programmatic travel in 2024 following our Summer Research in Mathematics visit in 2023. The second author was partially supported by the AMS-Simons Research Enhancement Grants for PUI Faculty while working on this project. The three authors also thank Dr. Lara Ismert for her conversations that inspired us to use Galois closed codes to move freely between the Hermitian and Euclidean inner products.

### v4

@@ 5 Galois Closed QC Codes and Their Duals
Let $R_{q^{2},m}^{\ell}$ be as in (25). Let $\mathcal{C}_{u}$ be the $m^{u}\ell$-length QC code whose constituent codes in  (27) and (28) are all Galois closed and whose minimum distances satisfy (10). Additionally, assume that the constituent codes in (28) are HSO. Then we have an infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of HSO $q^{2}$-ary QC Galois closed codes where $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,s}$ for each $u\geq 1$. Moreover,
  • (i)
    If $d(C_{1,t+1})\geq c\sqrt{\ell}$ for some constant $c$, then all codes in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound for their minimum distances.
  • (ii)
    Otherwise, there exist a positive integer $L>1$ so that the all codes in the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound for their minimum distances.
Remark 5.50.
Analogous results to those of Theorem 5.49 also hold when the constituent codes in (28) are HSD provided that $x^{m}-1$ has only x1x-1 as its self-reciprocal factor.
6 Constructions of QECCs from QC Codes
Given the construction of $\ell$-QC codes and results seen in cem; lingsole, along with the collection of quantum constructions in Subsection 2.4, we present in this section a construction of quantum codes with good parameters.
@@ 6 Constructions of QECCs from QC Codes
$R_{4,7}^{3}$ $\cong$ $\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{3}+x+1\rangle}\right)^{3}\oplus\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{3}+x^{2}+1\rangle}\right)^{3}\oplus\left(\frac{\mathbb{F}_{4}[x]}{\langle x-1\rangle}\right)^{3}$
@@
For any element $\gamma\in\mathbb{F}_{4}(\alpha)\setminus\mathbb{F}_{4}$, let $C_{1,1}^{\prime}=\left\langle(\gamma,\gamma,\gamma)\right\rangle\subset\mathbb{F}_{4}^{3}(\alpha)$ , $C_{1,1}^{\prime\prime}=(C_{1,1}^{\prime})^{\perp_{E}}\subset\mathbb{F}_{4}^{3}(\alpha^{3})$, and $C_{1,1}=\mathbb{F}_{4}^{3}$ be the constituent codes, whose parameters are $3,1,3]_{64}$, $3,2,2]_{64}$, and $3,3,1]_{4}$, respectively. Notice that $C_{1,1}^{\prime},C_{1,1}^{\prime\prime}$ are Galois closed by Proposition 5.45 and $C_{1,1}$ is trivially HDC and Galois closed. gray The minimum distance of the cyclic code associated to $\mathcal{C}$ and all its associated cyclic subcodes are described below
@@
WE NEED TO ADJUST THE REFERENCE BELOW TO JENSEN? There is a considerable difference between GO and Jensen bound for this case
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC code which, by Theorem 5.47, is also Galois-closed. By Lemma 2 and Theorem 2.8, $\mathcal{C}$ is a $21,12,7]_{4}-$HDC (and also EDC) QC code, where the minimum distance was computed using Magma. By Corollary 2.16, it is possible to obtain a $21,3,d\geq 7]]_{4}-$quantum code $Q$ from the QC code $\mathcal{C}$. According to table, the best-known possible minimum distance for a quaternary QECC with $n=21$ and $k=3$ is $7$.
@@
Ex. 6.51 and Lemma 2.17 - item (2) $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$ $\geq 6$
Code Table table 4-5 5-6 5-7 6-8 7-8
@@
Ex. 6.51 and Lemma 2.17 - item (1) $\geq 7$ $\geq 7$ $\geq 7$ $\geq 7$ $\geq 7$
Code Table table 7-10 8-11 8-11 9-11 9-12
Table 4: Parameters of good QECCs from our code construction
By Lemma 2.17, it is still possible to construct more good QECCs from the code $\mathcal{C}$ which are listed in Table 4. Although there is no significant improvement in the minimum distances for any of the above parameters, most of our codes achieve values that are within one or two of the best-known lower bounds. It is also worth noting that our example exhibits lower bounds for the minimum distance, so it may be possible to obtain codes with larger minimum distances.
@@
Still working with the decomposition of $R_{4,7}^{3}$ provided in Example 6.51, assume the same constituent codes $C_{1,1}^{\prime}$ and $C_{1,1}^{{}^{\prime\prime}}$ but let $C_{1,1}=\langle(0,1,\beta),(1,0,\beta^{2})\rangle$ be the last constituent code, where $\beta\in\mathbb{F}_{4}$ so that $\beta^{2}=\beta+1$. The code $C_{1,1}$ parameters $3,2,2]_{4}$ and is EDC.
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC code. By Corollary 3.26, we have an infinity family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of EDC QC codes whose their parameters are
$\left[3\cdot 7^{u},\frac{3\cdot 7^{u}+1}{2},\geq 2\cdot 7^{u-1}\right]_{4}.$}
From Corollary 2.16, it is possible to obtain the infinity family of quantum codes $\{Q_{u}\}_{u\geq 1}$ with parameters $\left[\left[3\cdot 7^{u},1,\geq 2\cdot 7^{u-1}\right]\right]_{4}$.
example Let $q=2$, $m=7$, $\ell=8$, and $\alpha$ be a primitive $7$-th root of unity. Assume the following decomposition of $R_{2,7}^{8}=\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$. Thus
@@
Let $C_{11}^{\prime}$ be a $8,3,6]_{8}$ MDS code (its construction can be checked in table), $C_{11}^{{}^{\prime\prime}}=C_{11}^{{}^{\prime}\perp_{E}}$ a $8,5,4]_{8}$ (also MDS) code, and $C_{11}$ be a EDC $8,6,2]_{2}-$code with basis
$B=\{$ $(1,0,0,1,0,0,0,0),(0,1,0,1,0,0,0,0),(0,0,1,1,0,0,0,0),(0,0,0,0,1,0,0,1),(0,0,0,0,0,1,0,1),$
$(0,0,0,0,0,0,1,1)\}$
ote that $d(D_{3})=7$, $d(D_{2,3})=3$, and $d(D_{1,2,3})=1$. So from Lemma 2 and Theorem 4, the code $\mathcal{C}_{1}$ is a $56,30,\geq 6]_{2}$ 88-QC code. The current best-known $56,30]_{2}$ code has minimum distance $10$. By Corollary 2.16, we obtain a quantum code $56,4,\geq 6]]_{2}$the best known minimum distance is $13$. AA: I feel that the phrases I colored in red directly above short-sell our results because our lower-bound for the minimum distance is far away from the best known. G.: I agree By Corollary 3.26, we obtain a $\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2}$ QC code. When $u=3$, $d\geq\sqrt{8\cdot 7^{u}}$, which satisfies the square-root bound. G. Gustavo´s suggestion ote that $d(D_{3})=7$, $d(D_{2,3})=3$, and $d(D_{1,2,3})=1$. So from Lemma 2 and Theorem 4, the code $\mathcal{C}_{1}$ is a $56,30,\geq 6]_{2}$ 88-QC code By Corollary 3.26, we obtain a $\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2}$ QC code. When $u=3$, $d\geq\sqrt{8\cdot 7^{u}}$, which satisfies the square-root bound. By Theorem 4.33, for $c=5$, we obtain $\left\lceil\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}\right\rceil+2=5$ and an infinite family $\{\mathcal{C}_{u}\}_{u\geq 5}$ with parameters
$\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2},$
where this bound to minimum distance is greater than $c\sqrt{n}$. magentaAA: I moved the last Example to the end of Section 4 as what’s impressive about it is that it satisfies the square-root and square-root-like lower bounds. As the classical code yields a family of quantum codes $\{\mathcal{Q}_{u}\}_{u\geq 2}$ with parameters
$\left[\left[8\cdot 7^{u},4,\geq 2\cdot 7^{u-1}\right]\right]_{2}.$
, I feel that it is not too impressive. as when $u=1$, we have a $\left[\left[56,4,2\right]\right]_{2}$ code, and the best known QECC for these pramaters has lower bound $13$. G.: Our first classical/quantum code is always poor due u1u-1, but it turns better from $u\geq 2$. By Corollary 2.16, we obtain a family of quantum codes $\{\mathcal{Q}_{u}\}_{u\geq 2}$ with parameters
$\left[\left[8\cdot 7^{u},4,\geq 2\cdot 7^{u-1}\right]\right]_{2}.$
7 Conclusion
In this work, we introduced a recursive construction of infinite families of quasi-cyclic codes whose parameters are determined by their constituent codes. By carefully selecting these constituents, we obtained families that are self-orthogonal, dual-containing, or self-dual with respect to the Euclidean and Hermitian inner products. A key feature of our construction is that both the dimension and a lower bound for the minimum distance can be explicitly determined from the parameters of the initial constituent codes, allowing us to study how the parameters evolve throughout the family. In particular, we showed that the resulting codes satisfy a square-root-like lower bound on the minimum distance which ensures a nontrivial distance growth as the code length increases. Moreover, we can use the CSS construction to establish the existence of quantum error-correcting codes with good parameters that arise from our families.
For future work, we aim to gain a deeper understanding of how using MDS codes as constituent codes affects the minimum distance of the respective QC Code. We also plan to explore sharper lower bounds for the minimum distance to improve the parameters of both the resulting QC code and its respective quantum code. Additionally, another plan is to relax the restriction in Section 3.2 that the constituent codes $C_{u,i}$ are given by $m^{u-1}$-copies of $C_{1,i}$. While this assumption simplified the analysis, allowing more general choices of $C_{u,i}$ may yield families with improved parameters. In each of these future directions, we aim to find the existence of quantum codes with good parameters.
@@ 7 Conclusion
n general, describing the parameters of infinity families of codes is not an easy task. In this paper, we have presented families of QC codes over $\mathbb{F}_{q}$, making explicit their dimension and a lower bound to their minimum distances. In addition, based on their constituent code decomposition, we also explored their dualities according to Euclidean and Hermitian inner products. We have also addressed when this lower bound to minimum distance is greater than the square-root-like bound, which there are only a few constructions in the literature achieving this goal. Finally, some quantum codes have been reported, which are comparable with those ones depicted on the Code Table table.
@@
AA: this paragraph needs to be changed as we are no longer use GO bound. For future work, we aim to gain a deeper understanding of how using MDS codes as constituent codes affects the minimum distance of the respective QC code. It is also important to explore whether the GO-bound can be improved, since doing so would lead to better lower-bound estimates for the minimum distances of both our QC codes and their corresponding quantum codes.
8 Acknowledgements
@@
The authors would like to thank the Institute for Computational and Experimental Research in Mathematics (ICERM) for providing funding and space for us to work in a mathematically rich environment during our week-long visit in Summer 2025. The first two authors would also like to thank Simons Laufer Mathematical Sciences Institute (SLMath) for providing funding for post-programmatic travel in 2024 following our Summer Research in Mathematics visit in 2023. The first author was supported by project FAPEMIG RED-00133-21. The second author was partially supported by the AMS-Simons Research Enhancement Grants for PUI Faculty while working on this project. The three authors also thank Dr. Lara Ismert for her conversations that inspired us to use Galois closed codes to move freely between the Hermitian and Euclidean inner products and Dr. Markus Grassl for reading the first version of this manuscript and pointing out some inconsistencies.
[…]